OpenStudy (anonymous):

2x+y=-1 -4x+6y=6 solve with substitution and elimination, show work please.

OpenStudy (anonymous):

can anyone help?

OpenStudy (anonymous):

2x+y=-1 ...................... (1) -4x+6y=6 ...................... (2) from (1) y = -1 -2x we substitute it in (2) we get -4x + 6( -1 -2x ) = 6 -4x -6 -12x = 6 -16x = 12 -8x = 6 -4x = 3 x = -3/4 x = -0.75

OpenStudy (anonymous):

Substitution: Rewrite the top equation as y=-1-2x and plug into the bottom equation: -4x + 6(-1-2x)=6 -4x -6 - 12x = 6 -16x = 12 x = 12/-16 x = - 3/4 plug this into one of the original equations y = -1-2(-3/4) y = -4/4 + 6/4 y = 2/4 y = 1/2 Elimination: It will be easier to multiply the top equation by 2 4x+2y=-2 Then add the second equation 4x - 4x + 2y + 6y = -2 + 6 8y = 4 y = 4/8 = 1/2 Plug this into one of the original equations: 2x + (1/2) = -1 2x = -2/2-1/2 2x = - 3/2 x = (-3/2)/2 x=-3/4

OpenStudy (anonymous):

S0fw0N, y should be positive.